A tiger running 100m race accelerates for one third of total time and then moves with uniform speed then find total time taken by tiger to run 100 m if acceleration of tiger is 8 m/s-2

Grade:11

3 Answers

Vikas TU
6 years ago
For the first t/3 sec, acceleration of tiger is 8 m/s^2
S1 = u1t1+1/2 at1^2
S1 = 0+ ½ x 8x (t/3)2
S1 = 4t2/9
Velocity = at1 = 8t/3;
S2 = velocity x t2
    = 8t/3 x 2t/3
    = 16t2/9
 
Total distance = 4t2/9 +16t2/9
100 = 20t2/9
900 = 20 t2
t =√45 sec.
Saksham sharma
3 years ago
Let total  time be t  and total distance be S              now for some distance .and the distance covered in time 1/3t is s1 so by using s=ut+1/2at² we get relation between s1 and time that is s1=4t²/9 and now acc . Is constant so  velocity is taken out by v=u+at we get that v=8t/3  and put this as u in s2 equation 2 of motion we get that s2=16t²/9 now w total distance S=s1+s2 hence putting value of all we get  time √45=3√5 sec
 
Dhyaan M
one year ago
An object moving with speed of 6.25m/s, is decelerated at the rate given by dv/dt = 2.5root v where v is the instantaneous speed. The time taken by the object to come to rest, would be